# The complete action for $\mathcal{N}=2$ de Sitter pure supergravity

### v1

@@ 1 Introduction
On the one hand, both the primordial and current Universe seem to be well approximated by de Sitter geometry. On the other hand, the non-renormalisability of Einstein’s theory of gravity calls for an extension of it that should behave better in the high-energy limit. Supergravities were born in the late seventies [1, 2] with the hope to give a renormalisable model of quantum field theory containing Einstein’s theory as a consistent truncation. This hope was later shown to be unrealistic, string theory offering more promises to attain UV finiteness while at the same time containing a massless spin-2 mode associated with the graviton, see e.g. [3]. More recently, however, the works [4, 5, 6, 7, 8, 9] and references therein prompted a renewed interest in supergravity theories. The very first paper where supergravity was considered in a de Sitter background is [10] by Pilch, van Nieuwenhuizen, and Sohnius. At that time, when supergravity theories were born, our Universe was believed to have no cosmological constant, and the paper [10] appeared to be of more academic relevance than phenomenological necessity. Still, this paper is seminal in that, for the first time, it clearly described the problems encountered in constructing supergravity theories in de Sitter (dS4) spacetime due to a clash between the reality conditions and the necessary gauge invariance to be imposed on a vector-spinor to only propagate the helicity degrees of freedom $\pm 3/2$. Supergravity favours a negative cosmological constant [11] — although this was not stressed in that seminal work that also borrowed a misleading terminology in use at the time, when the maximally symmetric spacetime of negative constant curvature was termed “de Sitter” instead of “anti-de Sitter” —, therefore it was crucial that the paper [10] pointed out the possible existence of a consistent supergravity theory in de Sitter spacetime at the price of a ghostly graviphoton, also showing that the minimal number of supersymmetries in dS4 supergravity is ${\cal N}=2\,$.
@@
Let us now review what is known for the two kinds of obstacles to the existence of unitary, unbroken de Sitter supersymmetry; see also [39] for a discussion of de Sitter $\mathcal{N}=1$ supergravity with spontaneously broken local supersymmetry, as well as the review given in [40].
@@ 3 Consistent interactions at cubic order
$2\,\sigma\,\alpha_{2}^{2}+\alpha_{1}^{2}=0\;,\quad\alpha_{1}=\frac{\kappa}{2}\;.$ (5)
@@ 4.1 Superalgebra, fields and gauge parameters
$\mathcal{R}_{\mu\nu}=$ $\,\mathcal{R}_{\mu\nu}^{a}(P)~P_{a}+\tfrac{1}{2}\mathcal{R}_{\mu\nu}{}^{ab}(M)M_{ab}+\bar{\mathcal{R}}_{\mu\nu}(Q)\,Q+\bar{Q}\,\mathcal{R}_{\mu\nu}(Q)+\mathcal{R}_{\mu\nu}(T)T\;.$ (14)
@@
where $R_{\mu\nu}{}^{ab}=2\partial_{[\mu}\omega_{\nu]}{}^{ab}+2\omega_{[\mu}{}^{cb}\omega_{\nu]}{}_{c}{}^{d}$ is the Riemann tensor. Also, $\mathcal{R}_{\mu\nu}(T)=F_{\mu\nu}+\bar{\psi}_{[\mu}i\gamma^{5}\psi_{\nu]}\;$.
@@ 4.2 The action
$+\frac{1}{2\sqrt{2}\kappa}\bar{\psi}_{\mu}\gamma^{5}(F^{\mu\nu}-i\gamma^{5}(*F)^{\mu\nu})\psi_{\nu}\Bigg)+\mathcal{O}(\psi^{4})\,$,
@@ Appendix A Gravitational coupling of the Dirac spin-3/2 field
The non-Abelian cubic interaction term involving a the massless spin-2 field $h_{\mu\nu}$ and a massless Dirac spin-3/2 field, $\phi_{\mu}$, is $S_{1}^{(h-\phi)}=\int d^{4}x\sqrt{-\bar{g}}\,\alpha^{(h-\phi)}L_{1}^{(h-\phi)}$, with
$L_{1}^{(h-\phi)}=$ $\;\bar{\phi}_{\alpha}\gamma^{\alpha\beta\gamma}\gamma^{\mu\nu}\gamma^{5}\phi_{\gamma}\nabla_{\mu}h_{\nu\beta}+2\bar{\phi}_{\delta}\gamma^{\gamma\delta\alpha}\gamma^{5}\nabla_{\gamma}\phi_{\alpha}h$ (22)
@@
$\delta^{(h-\phi)}_{1}h_{\mu\nu}$ $=\alpha^{(h-\phi)}\Bigl(\bar{\epsilon}\gamma_{(\mu}\gamma_{5}\phi_{\nu)}-\bar{\phi}_{(\mu}\gamma_{\nu)}\gamma_{5}\epsilon\Bigr)\,$, (23)
$\delta^{(h-\phi)}_{1}\phi_{\mu}$ $=\alpha^{(h-\phi)}\Bigl(\gamma^{\alpha\beta}\epsilon\nabla_{\alpha}h_{\beta\mu}-i\lambda\gamma^{\nu}\epsilon h_{\mu\nu}+2i\lambda\gamma_{\nu}\phi_{\mu}\xi^{\nu}-\gamma_{\alpha\beta}\phi_{\mu}\nabla^{\alpha}\xi^{\beta}+8\Phi_{\mu\nu}\xi^{\nu}\Bigr)\,$,
$\delta^{(h-\phi)}_{1}\bar{\phi}_{\mu}$ equals the Dirac conjugate of $\delta^{(h-\phi)}_{1}\phi_{\mu}$ , and $\delta^{(h-\phi)}_{1}A_{\mu}=0$ .
@@ Appendix B Electromagnetic coupling of the Dirac spin-3/2 field
The non-Abelian coupling between the massless spin-1 field and the Dirac spin-3/2 is $S^{(A-\phi)}_{1}=\int d^{4}x\sqrt{-\bar{g}}\,\alpha^{(A-\phi)}L^{(A-\phi)}_{1}$ , with
$L^{(A-\phi)}_{1}$=  $2\Bigl(\bar{\phi}_{\mu}\gamma^{5}(F^{\mu\nu}-i\gamma^{5}(*F)^{\mu\nu})\phi_{\nu}+i\lambda A_{\alpha}\bar{\phi}_{\gamma}\gamma^{\alpha\beta\gamma}\gamma^{5}\phi_{\beta}\Bigr)\;.$ (24)
@@
$\delta^{(A-\phi)}_{1}\phi_{\mu}$ $=\alpha^{(A-\phi)}\Bigl(-2i\lambda\phi_{\mu}\pi+2i\lambda A_{\mu}\epsilon-\frac{1}{2}F_{\alpha\beta}\gamma^{\alpha\beta}\gamma_{\mu}\epsilon\Bigr)\,$,
$\delta^{(A-\phi)}_{1}A_{\mu}$ $=\alpha^{(A-\phi)}2\sigma\Bigl(\bar{\epsilon}\gamma^{5}\phi_{\mu}-\bar{\phi}_{\mu}\gamma^{5}\epsilon\Bigr)\,$, (25)
$\delta^{(A-\phi)}_{1}\bar{\phi}_{\mu}$ equals the Dirac conjugate of $\delta^{(A-\phi)}_{1}\phi_{\mu}$ , and $\delta^{(A-\phi)}_{1}h_{\mu\nu}=0$ .

### v2

@@ 1 Introduction
On the one hand, both the primordial and current Universe seem to be well approximated by de Sitter geometry. On the other hand, the non-renormalisability of Einstein’s theory of gravity calls for an extension of it that should behave better in the high-energy limit. Supergravities were born in the late seventies [1, 2] with the hope to give a renormalisable model of quantum field theory containing Einstein’s theory as a consistent truncation. This hope was later shown to be unrealistic, string theory offering more promises to attain UV finiteness while at the same time containing a massless spin-2 mode associated with the graviton, see e.g. [3]. More recently, however, the works [4, 5, 6, 7, 8, 9] and references therein prompted a renewed interest in supergravity theories. The very first paper where supergravity was considered in a de Sitter background is [10] by Pilch, van Nieuwenhuizen, and Sohnius. At that time, when supergravity theories were born, our Universe was believed to have no cosmological constant, and the paper [10] appeared to be of more academic relevance than phenomenological necessity. Still, this paper is seminal in that, for the first time, it clearly described the problems encountered in constructing supergravity theories in de Sitter (dS4) spacetime due to a clash between the reality conditions and the necessary gauge invariance to be imposed on a vector-spinor to only propagate the helicity degrees of freedom $\pm 3/2$. Supergravity favours a negative cosmological constant [11] — although this was not stressed in the seminal work [11], which also adopted misleading terminology in use at the time, when the maximally symmetric spacetime of negative constant curvature was termed “de Sitter” instead of “anti-de Sitter” —, therefore it was crucial that the paper [10] pointed out the possible existence of a consistent supergravity theory in de Sitter spacetime at the price of a ghostly graviphoton, also showing that the minimal number of supersymmetries in dS4 supergravity is ${\cal N}=2\,$.
@@
Let us now review what is known for the two kinds of obstacles to the existence of unitary, unbroken de Sitter supersymmetry. [^1] For references about broken $\mathcal{N}=1$ supersymmetry and de Sitter, see e.g. [39, 40, 41, 42, 43, 44].
@@ 3 Consistent interactions at cubic order
$2\,\sigma\,\alpha_{2}^{2}+\alpha_{1}^{2}=0\;,\quad\alpha_{1}=-\frac{\kappa}{2}\;.$ (5)
@@ 4.1 Superalgebra, fields and gauge parameters
$\mathcal{R}_{\mu\nu}=$ $\,\mathcal{R}_{\mu\nu}^{a}(P)\,P_{a}+\tfrac{1}{2}\mathcal{R}_{\mu\nu}{}^{ab}(M)M_{ab}+\bar{\mathcal{R}}_{\mu\nu}(Q)\,Q+\bar{Q}\,\mathcal{R}_{\mu\nu}(Q)+\mathcal{R}_{\mu\nu}(T)T\;.$ (14)
@@
where $R_{\mu\nu}{}^{ab}=2\partial_{[\mu}\omega_{\nu]}{}^{ab}+2\omega_{[\mu}{}^{cb}\omega_{\nu]}{}_{c}{}^{a}$ is the Riemann tensor. Also, $\mathcal{R}_{\mu\nu}(T)=F_{\mu\nu}-\frac{1}{\sqrt{2}\kappa}\bar{\psi}_{[\mu}\gamma^{5}\psi_{\nu]}\;$.
@@ 4.2 The action
$-\frac{1}{2\sqrt{2}\kappa}\bar{\psi}_{\mu}\gamma^{5}(F^{\mu\nu}-i\gamma^{5}(*F)^{\mu\nu})\psi_{\nu}\Bigg)+\mathcal{O}(\psi^{4})\,$,
@@ Appendix A Gravitational coupling of the Dirac spin-3/2 field
The non-Abelian cubic interaction term involving a the massless spin-2 field $h_{\mu\nu}$ and a massless Dirac spin-3/2 field, $\phi_{\mu}$, is $S_{1}^{(\phi-h)}=\int d^{4}x\sqrt{-\bar{g}}\,\alpha_{1}L_{1}^{(\phi-h)}$, with
$L_{1}^{(\phi-h)}=$ $\;\bar{\phi}_{\alpha}\gamma^{\alpha\beta\gamma}\gamma^{\mu\nu}\gamma^{5}\phi_{\gamma}\nabla_{\mu}h_{\nu\beta}+2\bar{\phi}_{\delta}\gamma^{\gamma\delta\alpha}\gamma^{5}\nabla_{\gamma}\phi_{\alpha}h$ (22)
@@
$\delta^{(\phi-h)}_{1}h_{\mu\nu}$ $=\alpha_{1}\Bigl(\bar{\epsilon}\gamma_{(\mu}\gamma_{5}\phi_{\nu)}-\bar{\phi}_{(\mu}\gamma_{\nu)}\gamma_{5}\epsilon\Bigr)\,$, (23)
$\delta^{(\phi-h)}_{1}\phi_{\mu}$ $=\alpha_{1}\Bigl(\gamma^{\alpha\beta}\epsilon\,\nabla_{\alpha}h_{\beta\mu}-i\lambda\gamma^{\nu}\epsilon\,h_{\mu\nu}+2i\lambda\gamma_{\nu}\phi_{\mu}\xi^{\nu}-\gamma_{\alpha\beta}\phi_{\mu}\nabla^{\alpha}\xi^{\beta}+8\Phi_{\mu\nu}\xi^{\nu}\Bigr)\,$,
$\delta^{(\phi-h)}_{1}\bar{\phi}_{\mu}$ equals the Dirac conjugate of $\delta^{(\phi-h)}_{1}\phi_{\mu}$ , and $\delta^{(\phi-h)}_{1}A_{\mu}=0$ .
@@ Appendix B Electromagnetic coupling of the Dirac spin-3/2 field
The non-Abelian coupling between the massless spin-1 field and the Dirac spin-3/2 is $S^{(\phi-A)}_{1}=\int d^{4}x\sqrt{-\bar{g}}\,\alpha_{2}L^{(\phi-A)}_{1}$ , with
$L^{(\phi-A)}_{1}$=  $2\Bigl(\bar{\phi}_{\mu}\gamma^{5}(F^{\mu\nu}-i\gamma^{5}(*F)^{\mu\nu})\phi_{\nu}+i\lambda A_{\alpha}\bar{\phi}_{\gamma}\gamma^{\alpha\beta\gamma}\gamma^{5}\phi_{\beta}\Bigr)\;.$ (24)
@@
$\delta^{(\phi-A)}_{1}\phi_{\mu}$ $=\alpha_{2}\Bigl(-2i\lambda\phi_{\mu}\pi+2i\lambda A_{\mu}\epsilon-\frac{1}{2}F_{\alpha\beta}\gamma^{\alpha\beta}\gamma_{\mu}\epsilon\Bigr)\,$,
$\delta^{(\phi-A)}_{1}A_{\mu}$ $=2\,\alpha_{2}\,\sigma\Bigl(\bar{\epsilon}\gamma^{5}\phi_{\mu}-\bar{\phi}_{\mu}\gamma^{5}\epsilon\Bigr)\,$, (25)
$\delta^{(\phi-A)}_{1}\bar{\phi}_{\mu}$ equals the Dirac conjugate of $\delta^{(\phi-A)}_{1}\phi_{\mu}$ , and $\delta^{(\phi-A)}_{1}h_{\mu\nu}=0$ .